6+3 [2x-4(3x-2)]
help :(

Found 2 solutions by rfer, tommyt3rd:
Answer by rfer(16322)

6+3[2x-12x+8]
6+3[-10x+8]
6-30x+24=
-30x+30

Answer by tommyt3rd(5050)

6+3 [2x-4(3x-2)]=
6+3[2x-12x+8]=
6+3[8-10x]=
6+24-30x=
30-30x